# Periodic classical trajectories and quantum scars in many-spin systems

### v3

@@ VI Quantum spins
As discussed in the introduction, the present work was initially motivated by our interest to explore whether periodic orbits in classical spin lattices imply the existence of many-body scar eigenstates in quantum spin lattices. A quantum scar in a semiclassical limit can be associated with self-interference of a quantum wavepacket propagating along a periodic classical trajectory and thereby forming an atypical “standing wave”[115]. In a many-dimensional phase space, the presence of many Lyapunov instabilities along that trajectory is expected to make the above wavepacket too broad before it could start interfering with itself, thereby reducing the chances for a distinct scar eigenstate to emerge. From such a perspective, the weaker the Lyapunov instabilities, the better are the chances for observing quantum scars. In this regard, our classical investigation in Section IV indicates that the chances of observing the scar effects in many-spin systems are, actually, somewhat better than what one could expect from the knowledge of the Lyapunov spectra of ergodic trajectories[126]. The reason is that one can choose the parameters of the Hamiltonian such that the largest Lyapunov exponent for a periodic trajectory is significantly smaller than that for an ergodic trajectory (cf. Figs. 3 and 24), and, moreover, there exist rather long spin chains [see Figs. 4(b,c)] for which the Lyapunov instabilities in the vicinity of the periodic trajectories are completely suppressed.
@@ VI.1 Suppressed initial thermalization
where $|\text{inf}\rangle_{L-1}$ is a pure state randomly sampled in the Hilbert space of spins on sites from 2 to $L$  [^1] A pure quantum state randomly sampling the infinite temperature equilibrium in a Hilbert space of dimension $N\gg 1$ is obtained as $|\text{inf}\rangle=\frac{1}{\sqrt{2^{L-1}}}\sum^{2^{L-1}}_{n=0}c_{n}|b_{n}\rangle$, where $|b_{n}\rangle$ are basis states, and $c_{n}=|c_{n}|e^{i\phi_{n}}$ are complex amplitudes with phases $\phi_{n}$ chosen randomly from the interval $[0,2\pi)$ and absolute values $|c_{n}|$ chosen using the probability distribution $P(|c_{n}|^{2})=Ne^{-N|c_{n}|^{2}}$ (see Refs.[162, 163]). Quantum typicality [77, 162, 163] implies that the energy expectation value for $|\Psi^{\text{inf}}\rangle$ is exponentially close to the value for the infinite temperature canonical ensemble, which is, in turn, equal to zero [see Eq.(III)]. Since the energy expectation value for $|\Psi^{\text{up}}\rangle$ is also zero, both $|\Psi^{\text{up}}\rangle$ and $|\Psi^{\text{inf}}\rangle$ belong to the same “energy shell” in the sense of the canonical ensemble.
@@
We, finally, note that Fig. 16 also includes the results of classical simulations for an ensemble of initial conditions devised to imitate the relaxation of quantum spins $S$ starting from the initial state $|\Psi^{\text{up}}\rangle$. In the simulations, each classical spin has length 1. In order to represent a quantum state $|m=S\rangle$, the range $[-1,1]$ of the $z$-projection of a classical spin is divided into $12S+$ 1 equal intervals of length $\frac{2}{2S+1}$ and then the classical projection $S_{z}$ is randomly sampled from the topmost interval $[1-\frac{2}{2S+1},1]$, while the $x$- and $y$-projections are chosen as $S_{x}=\sqrt{1-S_{z}^{2}}\cos\varphi$ and $S_{y}=\sqrt{1-S_{z}^{2}}\sin\varphi$, where $\varphi$ is a random angle from the interval $[0,2\pi)$. On the one hand, the above ensemble becomes increasingly close to the considered periodic classical trajectories as the value of quantum spin $S$ increases. On the other hand, as one can see in Fig. 16, the average for that ensemble exhibits an excellent agreement with the quantum simulations starting already with $S=1$. This is yet another indication, in addition to the observations of Ref.[68], that classical spin ensembles can quantitatively represent quantum spin dynamics for translationally invariant chains of spins $S\geq 1$ at high temperatures. The classicalization of higher spins was also recently reported at the operator level in [164], further highlighting the difference in dynamical properties between spin-$1/2$ and higher spins. In the present context, the success of the above simulation indicates the proximity of quantum dynamics to the semiclassical limit, which will be put forward below as a condition for the observability of distinct many-body quantum scars.
@@ VI.3 Quantum separatrix
In Fig. 21, we plot numerically computed $\mathcal{P}(\tilde{J})$ for different $S$. These plots show that, for $S=3/2$ and $S=2$, $\mathcal{P}(\tilde{J})$ exhibits well-defined peaks located with good precision at the classical separatrix value $\tilde{J}=J^{*}$. For $S=1/2$, $\mathcal{P}(J)$ does not show any clear peak, while, for $S=1$, the peak around $\tilde{J}=J^{*}$ does exist, but its maximum is not well defined. The above findings thus consistent with the clear evidence of quantum scars presented in Fig. 18 for the simulated spin-3/2 and spin-2 chains, the weaker evidence for spins 1 and the lack of the evidence for spins-1/2.
@@ VII Summary and discussion
Note added: After submitting the present work for publication, we were alerted by the authors of Refs.[171, 172, 173] about their closely related publications on the relations between classical periodic orbits and quantum many-body scars.
@@ Appendix B Upper bound on the largest Lyapunov exponent for classical spin dynamics
[…]
@@ Appendix F Functional form of $\lambda_{\text{p}}(J)$ near the separatrix $J=J^{*}$
[…]
@@
Since the separatrix has an unstable fixed point $\mathbf{S}^{*}$ characterized by the pair of Lyapunov exponents ±$\pm\lambda_{S}$, the period of a periodic trajectory in the vicinity of the separatrix logarithmically diverges as $J$ approaches $J^{*}$. Here the Lyapunov exponents ±$\pm\lambda_{S}$ are relevant, not because we investigate the instability of the fixed point, but because the reference periodic trajectory itself has a stretch that exhibits nearly exponential slowdown followed by the nearly exponential acceleration in the vicinity of the fixed point.
@@
$-f\Delta s_{1}^{2}+g\Delta s_{2}^{2}=\Delta J,$ (51)
@@
$=$ $\ C_{1}-\frac{1}{2\lambda_{S}}\log|\Delta J|\ \xrightarrow[|\Delta J|\to 0]{}\ -\frac{1}{2\lambda_{S}}\log|\Delta J|,$
@@
The time interval $T_{S}$ of the Lyapunov growth of $|\Delta\mathbf{S}(t)|$ is then followed by the time interval $T_{0}$ of large deviations from the fixed point; that interval is approximately independent of small $|\Delta J|$, because $T_{0}(J)\approx T_{0}(J^{*})+\left.\frac{dT_{0}}{dJ}\right|_{J=J^{*}}\Delta J\approx T_{0}(J^{*})$. At the end of $T_{0}$, $|\Delta\mathbf{S}(t)|$ crosses again the threshold value $|\Delta\mathbf{S}|_{\lambda}$ and starts approaching the fixed point in the Lyapunov regime controlled by the negative Lyapunov exponent $-\lambda_{S}$. Due to the time-reversal symmetry of the Hamiltonian dynamics, it takes another time $T_{S}$ until the trajectory arrives to the point of the closest approach to the fixed point, thereby completing the period $T$. The period $T$ can now be represented as:

### v4

@@ VI Quantum spins
As discussed in the introduction, the present work was initially motivated by our interest to explore whether periodic orbits in classical spin lattices imply the existence of many-body scar eigenstates in quantum spin lattices. A quantum scar in a semiclassical limit can be associated with self-interference of a quantum wavepacket propagating along a periodic classical trajectory and thereby forming an atypical “standing wave”[115]. In a many-dimensional phase space, the presence of many Lyapunov instabilities along that trajectory is expected[116] to make the above wavepacket too broad before it could start interfering with itself, thereby reducing the chances for a distinct scar eigenstate to emerge. From such a perspective, the weaker the Lyapunov instabilities, the better are the chances for observing quantum scars. In this regard, our classical investigation in Section IV indicates that the chances of observing the scar effects in many-spin systems are, actually, somewhat better than what one could expect from the knowledge of the Lyapunov spectra of ergodic trajectories[127]. The reason is that one can choose the parameters of the Hamiltonian such that the largest Lyapunov exponent for a periodic trajectory is significantly smaller than that for an ergodic trajectory (cf. Figs. 3 and 24), and, moreover, there exist rather long spin chains [see Figs. 4(b,c)] for which the Lyapunov instabilities in the vicinity of the periodic trajectories are completely suppressed.
@@ VI.1 Suppressed initial thermalization
where $|\text{inf}\rangle_{L-1}$ is a pure state randomly sampled in the Hilbert space of spins on sites from 2 to $L$  [^1] A pure quantum state randomly sampling the infinite temperature equilibrium in a Hilbert space of dimension $N\gg 1$ is obtained as $|\text{inf}\rangle=\frac{1}{\sqrt{2^{L-1}}}\sum^{2^{L-1}}_{n=0}c_{n}|b_{n}\rangle$, where $|b_{n}\rangle$ are basis states, and $c_{n}=|c_{n}|e^{i\phi_{n}}$ are complex amplitudes with phases $\phi_{n}$ chosen randomly from the interval $0,2\pi)$ and absolute values $|c_{n}|$ chosen using the probability distribution $P(|c_{n}|^{2})=Ne^{-N|c_{n}|^{2}}$ (see Refs.[163, 164]). Quantum typicality [77, 163, 164] implies that the energy expectation value for $|\Psi^{\text{inf}}\rangle$ is exponentially close to the value for the infinite temperature canonical ensemble, which is, in turn, equal to zero [see Eq.(III)]. Since the energy expectation value for $|\Psi^{\text{up}}\rangle$ is also zero, both $|\Psi^{\text{up}}\rangle$ and $|\Psi^{\text{inf}}\rangle$ belong to the same “energy shell” in the sense of the canonical ensemble.
@@
We, finally, note that Fig. 16 also includes the results of classical simulations for an ensemble of initial conditions devised to imitate the relaxation of quantum spins $S$ starting from the initial state $|\Psi^{\text{up}}\rangle$. In the simulations, each classical spin has length 1. In order to represent a quantum state $|m=S\rangle$, the range [1,1][-1,1] of the $z$-projection of a classical spin is divided into $2S+1$ equal intervals of length $\frac{2}{2S+1}$ and then the classical projection $S_{z}$ is randomly sampled from the topmost interval $1-\frac{2}{2S+1},1]$, while the $x$- and $y$-projections are chosen as $S_{x}=\sqrt{1-S_{z}^{2}}\cos\varphi$ and $S_{y}=\sqrt{1-S_{z}^{2}}\sin\varphi$, where $\varphi$ is a random angle from the interval $0,2\pi)$. On the one hand, the above ensemble becomes increasingly close to the considered periodic classical trajectories as the value of quantum spin $S$ increases. On the other hand, as one can see in Fig. 16, the average for that ensemble exhibits an excellent agreement with the quantum simulations starting already with $S=1$. This is yet another indication, in addition to the observations of Ref.[68], that classical spin ensembles can quantitatively represent quantum spin dynamics for translationally invariant chains of spins $S\geq 1$ at high temperatures. The classicalization of higher spins was also recently reported at the operator level in [165], further highlighting the difference in dynamical properties between spin-$1/2$ and higher spins. In the present context, the success of the above simulation indicates the proximity of quantum dynamics to the semiclassical limit, which will be put forward below as a condition for the observability of distinct many-body quantum scars.
@@ VI.3 Quantum separatrix
[…]
@@ VII Summary and discussion
Note added: After submitting the present work for publication, we were alerted by the authors of Refs.[172, 173, 174] about their closely related publications on the relations between classical periodic orbits in spin systems and quantum many-body scars. We share the optimism of the authors of Ref. [174] that the quantum interference effect associated with the periodic classical orbits can lead to “genuine” many-body quantum scars.
@@ Appendix B Upper bound on the largest Lyapunov exponent for classical spin dynamics
$\lambda_{\text{max}}=\frac{1}{T}\int_{0}^{T}\lambda_{\text{inst}}(t)dt,$ (33)
@@ Appendix F Functional form of $\lambda_{\text{p}}(J)$ near the separatrix $J=J^{*}$
$\lambda_{\text{p}}=\frac{1}{T}\int_{0}^{T}\lambda_{\text{inst}}(t)dt,$ (49)
@@
Since the separatrix has an unstable fixed point $\mathbf{S}^{*}$ characterized by the pair of Lyapunov exponents $\pm\lambda_{S}$, the period of a periodic trajectory in the vicinity of the separatrix logarithmically diverges as $J$ approaches $J^{*}$. Here the Lyapunov exponents $\pm\lambda_{S}$ are relevant, not because we investigate the instability of the fixed point, but because the reference periodic trajectory itself has a stretch that exhibits nearly exponential slowdown followed by the nearly exponential acceleration in the vicinity of the fixed point.
@@
$f\Delta s_{1}^{2}+g\Delta s_{2}^{2}=\Delta J,$ (51)
@@
$=$ $\ C_{1}-\frac{1}{2\lambda_{S}}\log|\Delta J|\ \xrightarrow[|\Delta J|\to 0]{}\ -\frac{1}{2\lambda_{S}}\log|\Delta J|,\ \ \ \ \$
@@
The time interval $T_{S}$ of the Lyapunov growth of $|\Delta\mathbf{S}(t)|$ is then followed by the time interval $T_{0}$ of large deviations from the fixed point; that interval is approximately independent of small $|\Delta J|$, because $T_{0}(J)\approx T_{0}(J^{*})+\left.\frac{dT_{0}}{dJ}\right|_{J=J^{*}}\Delta J\approx T_{0}(J^{*})$. At the end of $T_{0}$, $|\Delta\mathbf{S}(t)|$ crosses again the threshold value $|\Delta\mathbf{S}|_{\lambda}$ and starts approaching the fixed point in the Lyapunov regime controlled by the negative Lyapunov exponent $\lambda_{S}$. Due to the time-reversal symmetry of the Hamiltonian dynamics, it takes another time $T_{S}$ until the trajectory arrives to the point of the closest approach to the fixed point, thereby completing the period $T$. The period $T$ can now be represented as: